OpenStudy (anonymous):

Please help me! Solve for x: |4x + 12| = 16 x = 7, x = –7 x = 1, x = –1 x = 1, x = –7 x = –1, x = 7 Which answer is it??? Need help!

DivineSolar (divinesolar):

Lets walk through this.

OpenStudy (anonymous):

OK!

DivineSolar (divinesolar):

Subtract 12 from both sides 4x + 12 - 12 = 16 - 12 > This gives us what?

OpenStudy (anonymous):

4?

DivineSolar (divinesolar):

Correct 4x = 4 so now divide both sides by 4 > \[\frac{ 4x }{ 4 } = \frac{ 4 }{ 4 }\]

DivineSolar (divinesolar):

What do you get?

OpenStudy (anonymous):

1?

DivineSolar (divinesolar):

Correct so now we have x = ?

OpenStudy (anonymous):

1?

DivineSolar (divinesolar):

Yes x = 1, Now lets solve for our second x

OpenStudy (anonymous):

OK

DivineSolar (divinesolar):

So subtract 12 from both sides again 4x + 12 - 12 = -16 - 2 > gives us what?

OpenStudy (anonymous):

0?

DivineSolar (divinesolar):

No try this time instead of guessing :).

DivineSolar (divinesolar):

-12

DivineSolar (divinesolar):

- 16 - 12

OpenStudy (anonymous):

OK well I see 12 on the other side, So that's why I though 12-12 0. But Im not sure what to do really

DivineSolar (divinesolar):

-16 (negative minus a negative basically means your adding the 2 but they end up a negative so -2 - 2 is basically 2 +2 for negatives, so we would get -4)

DivineSolar (divinesolar):

So try again > 4x + 12 - 12 = -16 - 12

OpenStudy (anonymous):

OK I am following ..

OpenStudy (anonymous):

-7

DivineSolar (divinesolar):

I see you went ahead and did the division, correct. x = -7

OpenStudy (anonymous):

Oh well thank you SOOO much!!!

DivineSolar (divinesolar):

Sure thing.